OpenStudy (anonymous):

how to verify frac(sec theta/ csc theta- cot theta -frac(sectheta/csc theta +cot theta = 2csc theta

Nnesha (nnesha):

\[\large\rm \frac{ \sec \theta }{ \csc \theta -\cot \theta } -\frac{ \sec \theta }{ \csc \theta +\cot \theta }=2\csc \theta \] like this ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Can you please help

Nnesha (nnesha):

ye gimme a sec let me do it i'll try to find easy way

OpenStudy (anonymous):

thanks

Nnesha (nnesha):

okay alright so first we should write sec and csc in terms of sin or cos csc = ?? sec equal what ? do you know ?

OpenStudy (anonymous):

1/sin - csc and sec =1/cos

OpenStudy (anonymous):

i meant csc=1/sin

Nnesha (nnesha):

yes right first deal with the denominator\[\large\rm \frac{ \sec \theta }{ \color{ReD}{\csc \theta -\cot \theta }} -\frac{ \sec \theta }{\color{reD}{ \csc \theta +\cot \theta }}=2\csc \theta \] \[\huge\rm \frac{ \sec \theta }{ \frac{ 1 }{ \sin }-\frac{ \cos }{ \sin}} - \frac{ \sec \theta }{ \frac{ 1 }{ \csc }+\frac{ \cos }{ \sin } }\] cot =cos over sin

OpenStudy (anonymous):

Yes i got that far then i get stuck

Nnesha (nnesha):

\[\huge\rm \frac{ \sec \theta }{ \color{red}{\frac{ 1 }{ \sin }-\frac{ \cos }{ \sin}}} - \frac{ \sec \theta }{\color{red}{ \frac{ 1 }{ \csc }+\frac{ \cos }{ \sin } }}\] find common denominator of red part

OpenStudy (anonymous):

but i get 1/sin +cos/sin

OpenStudy (anonymous):

1-cos/sin - 1+cos/sin

Nnesha (nnesha):

if you have some work show it to me so i can find ur mistakes instead stating ovr

OpenStudy (anonymous):

1-cos/ sin -1+cos/sin sin/sin-sin/sin which equals 1 then i flip it to multiuply with 1/cos-1/cos

Nnesha (nnesha):

ugh gawd

Nnesha (nnesha):

brb i should refresh the page its lagging

OpenStudy (anonymous):

here is where i get lost

Nnesha (nnesha):

\[\large\rm \frac{ \sec }{ \frac{ 1+\cos }{ \sin } }-\frac{ \sec }{ \frac{ 1+\cos }{ \sin } }\]now multiply top with the reciprocal of the bottom fraction (change division to multiplication ) like for first one \[\large\rm \sec \times \frac{ \sin }{ 1-\cos } - \sec \times \frac{ \sin }{ 1+\cos }\]

OpenStudy (anonymous):

so then I get 1-cos(sec) /sin - 1+cos(sec)/sin

Nnesha (nnesha):

yes right we can change sec to 1 over cos \[\frac{ 1 }{ \cos } \times \frac{ \sin }{ 1-\cos } - \frac{ 1 }{ \cos } \times \frac{\sin }{ 1+\cos }\] \[\frac{ \sin }{ \cos(1-\cos) } -\frac{ \sin }{ \cos (1+\cos) }\] now find the common denominator

OpenStudy (anonymous):

cos (1-cos)(1+cos)

Nnesha (nnesha):

yes right \[\huge\rm \frac{ \sin\color{reD}{(1+\cos )}-\sin\color{reD}{(1-\cos)} }{ \cos (1-\cos)(1+\cos) }\] multiply sin with the denominator of 2nd fraction and multiply numerator of 2nd fraction with the denominator of 1st fraction that's how i got the red part now distribute parentheses by sin at the top and foil(1-cos)(1+cos)

OpenStudy (anonymous):

sin+sin cos - sin-sin cos

OpenStudy (anonymous):

then distribute? sorry wrong

Nnesha (nnesha):

sign mistake -sin times -cos = ?

Nnesha (nnesha):

yes foil (1-cos)(1+cos) which is at the denominator

OpenStudy (anonymous):

sin+sin cos

OpenStudy (anonymous):

cos (cos^2)

Nnesha (nnesha):

yes right \[\huge\rm \frac{ \sin +sincos-\sin +sincos }{ \cos(1-\cos)(1+\cos) }\] now (1-cos)(1+cos ) = ?

Nnesha (nnesha):

now it's not just cos ^2  what do you get when you multiply first term by 1st term of 2nd parentheses